OpenStudy (anonymous):

Help with 2 statistics problems: A seller of a machine claims that the machine cuts lumber to a mean length of 96in with a standard deviation of 0.35in. Assume the lengths are normally distributed and assume that the seller is correct. a. What is the probability that a single randomly selected peice of lumber would have a length greater than 96.15inches? b. What is the probability that the average length of 49 randomly selected pieces of lumber would be greater than 96.15in ? I don't need help solving the problems, just setting them up.

OpenStudy (anonymous):

For A I have started by: P(Y>96.15) = P(Z> [96.15-96]/0.35) = 1-0.6664 = 0.3336 I'm not sure if I have set this up correctly.

OpenStudy (anonymous):

you are trying to find a zscore for 96.15 from the mean ... right?

OpenStudy (perl):

the z score should be z= 0.42857

OpenStudy (anonymous):

im wondering if the minus is due to a left tailed table ...

OpenStudy (perl):

yes because her table is left tailed

OpenStudy (perl):

P(Z>.42857) = 1 - P( Z < .42857)

OpenStudy (perl):

I use a calculator and got the exact answer normalcdf ( 96.15,1E99, 96, .35) =0.3341176, which is 0.334 to three decimals

OpenStudy (anonymous):

same here :)

OpenStudy (anonymous):

Is this for part A? What is confusing to me is that one is asking for a single sample and the other is asking for the entire mean. . . I'm not sure what the different formulas are

OpenStudy (anonymous):

divide the sd by sqrt(n)

OpenStudy (anonymous):

\[z=\frac{x-\bar x}{\sigma/\sqrt{n}}\]

OpenStudy (anonymous):

in other words, your z score changes by a multiple of \(\sqrt{49}=7\)

OpenStudy (anonymous):

AH! I see now, thank you!

OpenStudy (anonymous):

youre welcome :)

OpenStudy (perl):

the new z score for the 49 lumber pieces is z = 3

OpenStudy (perl):

P(Z>3) = .00134996